OpenStudy (anonymous):

What is the slope of the line 2x + 5y = -10?

OpenStudy (raden):

ax + by = c has the slope m = -a/b

OpenStudy (anonymous):

...

OpenStudy (raden):

2x + 5y = -10 here known a=2, b=5 use the formula above

OpenStudy (anonymous):

you could just put it in y = mx + b form and m is the slope 2x + 5y = 10 5y = -2x + 10 y = -2/5x + 2 slope is -2/5